# Gaussian-LIC: Real-Time Photo-Realistic SLAM with Gaussian Splatting and LiDAR-Inertial-Camera Fusion

### v2

@@ I INTRODUCTION
Simultaneous localization and mapping (SLAM) is a pivotal technique in mixed reality and robotics. In recent years, emerging radiance field technologies such as Neural Radiance Fields (NeRF) [mildenhall2021nerf] and 3D Gaussian Splatting (3DGS) [kerbl20233d] have carved out a new direction in the realm of SLAM, namely radiance-field-based SLAM [tosi2024nerfs], aiming to simultaneously localize accurately and construct photo-realistic 3D maps. This holds significant importance for enhancing the capabilities of robots to model, understand, and interact with natural environments in real time [zheng2024gaussiangrasper, jin2024gs, chen2024splat, zuo2024fmgs].
Compared to NeRF which involves computationally intensive volumetric rendering based on ray tracing, 3DGS offers fast rendering speed with superior visual quality, showcasing a greater potential in real-time SLAM applications. Yet there still remain several issues for 3DGS-based SLAM. First, similar to existing NeRF-based SLAM [sucar2021imap, zhu2022nice, yang2022vox, wang2023co, johari2023eslam, sandstrom2023point, xin2024hero, rosinol2023nerf, naumann2023nerf], most 3DGS-based SLAM [yan2024gs, keetha2024splatam, yugay2023gaussian, ha2024rgbd, hu2024cg, peng2024rtg, sarikamis2024ig, zhu2024mgs, huang2024photo, matsuki2024gaussian, li2024sgs, zhu2024semgauss, ji2024neds] focus mainly on indoor scenarios using RGB-D or RGB sensors. When it comes to unbounded outdoor scenes or encountering challenging conditions such as violent ego-motion, varying illumination, and lack of visual textures, the performance of existing 3DGS-based SLAM might dramatically diminish. Second, for the outdoor scenarios, the modeling of the sky and varying camera exposure becomes important. Ignoring them can incur significant artifacts to 3DGS in unbounded scenes. Third, real-time performance is still a bottleneck. Constraining the number of or the dimension of attributes of 3D Gaussian representation [keetha2024splatam, peng2024rtg] could be effective indoors but might lead to deteriorated performance in unbounded outdoor scenarios.
@@ II-A Photo-Realistic RGB-D or RGB SLAM
Given sequential RGB-D or RGB inputs, abundant works attain brilliant results of localization and photo-realistic mapping in indoor environments. iMAP [sucar2021imap], the first NeRF-based SLAM that employs the implicit neural representation to achieve watertight online reconstruction, has pioneered a new era in SLAM. As a follow-up, NICE-SLAM [zhu2022nice] combines MLPs with hierarchical feature grids, markedly improving the quality of scene representation and realizing outstanding performance in larger indoor rooms. Further, Vox-Fusion [yang2022vox] utilizes octree to dynamically expand the volumetric neural implicit map, eliminating the need for pre-allocated grids. Adopting hash-grids, tri-planes, and neural point clouds as implicit neural representations respectively, Co-SLAM [wang2023co], ESLAM [johari2023eslam], and Point-SLAM [sandstrom2023point] get enhancement in both localization and reconstruction. Meanwhile, HERO-SLAM [xin2024hero] improves the robustness in the optimization of the neural map. NeRF employs time-consuming ray-based volume rendering, while 3DGS utilizes fast point-based rasterization, accelerating image view synthesis and producing promising rendering quality. Notably, GS-SLAM [yan2024gs], SplaTAM [keetha2024splatam], and Gaussian-SLAM [yugay2023gaussian] detailedly elucidate the significant advantages of 3DGS over existing map representations in SLAM tasks for online photo-realistic mapping. SGS-SLAM [li2024sgs], SemGauss-SLAM [zhu2024semgauss], and NEDS-SLAM [ji2024neds] integrate semantic feature embedding into Gaussians and demonstrate superb capabilities in online dense semantic mapping. Fusing Generalized Iterative Closest Point (G-ICP) and 3DGS, GS-ICP-SLAM [ha2024rgbd] shares covariances of Gaussians between tracking and mapping to minimize redundant computations. CG-SLAM [hu2024cg] additionally incorporates a depth uncertainty model to select valuable Gaussians and thereby improves tracking performance. By forcing each Gaussian to be either opaque or nearly transparent, RTG-SLAM [peng2024rtg] achieves real-time performance indoors with compact scene representation.
Several works have tried to operate solely on monocular images. Among them, NeRF-SLAM [rosinol2023nerf] and IG-SLAM [sarikamis2024ig] utilize the dense depth maps estimated from the tracking front-end DROID-SLAM [teed2021droid] as additional information to supervise the training of Instant-NGP [muller2022instant] and 3DGS. Correspondingly, NeRF-VO [naumann2023nerf] and MGS-SLAM [zhu2024mgs] instead employ the sparse visual odometry DPVO [teed2024deep] as a faster front-end with network-predicted dense depth for supervision. Concurrently, Photo-SLAM [huang2024photo] utilizes the classical visual odometry ORB-SLAM3 [campos2021orb] for accurate pose estimation and reconstructs a hybrid Gaussian map with ORB features. MonoGS [matsuki2024gaussian] introduces geometric regularization to address ambiguities in incremental reconstruction.
@@ II-B Photo-Realistic Multimodal SLAM
RGB-D and RGB approaches can realize fancy results in well-lit indoor scenes, but they would find it difficult to scale up to complex unbounded outdoor environments or cope with violent motions, lighting changes, and texture missing. Also, RGB-D sensors are often unavailable outdoors. To this end, fusing multimodal sensors such as LiDAR and IMU will be necessary and effective, which has been extensively proven in classical SLAM methods [zuo2019lic, zuo2020lic, shan2021lvi, lin2022r, zheng2022fast, lv2023continuous, lang2023coco] characterized by robustness and accuracy. Nevertheless, their primary focus lies in geometric mapping, resulting in geometrically precise yet visually simple representations of the scene. URF [rematas2022urban], DrivingGaussian [zhou2024drivinggaussian], StreetGaussians [yan2024street], PVG [chen2023periodic], LIV-GaussMap [hong2024liv], and LetsGo [cui2024letsgo] perform enhanced photorealistic reconstruction based on LiDAR-camera fusion, but all of them are offline methods taking a batch input.
Oriented to SLAM tasks, MM-Gaussian [wu2024mm] leverages accurate geometric structure information from a solid-state LiDAR to substitute the depth of the RGB-D, thus scaling up to the outdoor scenes. Specifically, it first performs a pure LiDAR odometry [vizzo2023kiss] for pose estimations and further refines the pose via color and depth rendering loss. Similar to SplaTAM [keetha2024splatam], it optimizes isotropic Gaussians without view-dependent effect initialized from LiDAR points. MM3DGS-SLAM [sun2024mm3dgs] fuses RGB-D camera and IMU measurements in a loosely-coupled manner, to enable more accurate and scale-aware pose estimation. Both of the aforementioned multimodal SLAM approaches are capable of functioning in outdoor environments. However, their mapping performance is suboptimal due to the lack of sky modeling and the challenges posed by varying camera exposure in outdoor settings. Additionally, they are not specially optimized for real-time applications.
@@ III-A Preliminary: 3D Gaussian Splatting
To produce a richly detailed dense map for fast and high-quality rendering, we represent the scene as a collection of anisotropic 3D Gaussians [kerbl20233d], each of which contains position $\bm{\mu}\in\mathbb{R}^{3}$ in the world coordinate, scale $\mathbf{S}\in\mathbb{R}^{3}$, rotation $\mathbf{R}\in\mathbb{R}^{3\times 3}$, opacity $o\in\mathbb{R}$, and three-degree spherical harmonics $\mathbf{SH}\in\mathbb{R}^{48}$ capturing the view-dependent appearance of the scene. Representing the Gaussian’s ellipsoidal shape, the covariance of each Gaussian is parameterized as:
@@
where $\mathbf{e}_{i}$ is a $3\times 1$ vector with its $i$-th element to be 1 and the other elements to be 0. Then ${\mathbf{e}_{3}^{\top}\hat{\bm{\mu}}}$ is the depth $d$ of the Gaussian in the camera frame. $\pi_{c}(\cdot)$ denotes the projection operation which transforms a point on the normalized image plane to a pixel. $\mathbf{J}$ is the Jacobian of the affine approximation of the projective transformation and $[\cdot]_{2\times 2}$ skips the third row and column to acquire a $2\times 2$ matrix. The projected 2D Gaussian affects the pixel $\bm{\rho}=\begin{bmatrix}u&v\end{bmatrix}^{\top}$ by the weight:
@@ III-B Multimodal Odometry and Mapping Data Preparation
Given data from multimodal sensors from an RGB camera, a 3D LiDAR, and an IMU, we adopt our previous work Coco-LIC [lang2023coco], a continuous-time odometry to tightly fuse the sensor data, delivering robust, real-time, and accurate pose estimation even in adverse scenarios. It utilizes a factor graph to incrementally optimize the continuous-time trajectory for every 0.1 seconds incorporating point-to-map LiDAR factors, frame-to-map visual factors, and inertial factors.
For photo-realistic reconstruction, 3D Gaussians in the map need to be bootstrapped by a set of 3D points. The better initialization leads to faster convergence and higher-quality rendering. Exactly, the LiDAR points provide handy and more precise geometric priors for Gaussian initialization [zhou2024drivinggaussian]. Our system supports both solid-state and mechanical spinning LiDARs. Specifically, after finishing the window optimization of the trajectory segment in the latest 0.1 seconds, we first downsample the LiDAR points in the time interval by retaining one out of every $N_{l}$ points. Then we colorize the remained LiDAR points by projection onto the latest image and remove those outside the corresponding camera FoV (field of view). Afterward, these colorized LiDAR points in the world frame are sent to the mapping thread, together with the latest image with the estimated camera pose, regarded as a frame.
Though offering accurate geometric priors, the region scanned by the LiDAR is limited compared to the camera, especially for the repetitive mechanical spinning LiDAR. Thus, we propose maintaining a visual sliding window similar to VINS-Mono [qin2018vins] to compute triangulated SFM points for compensation. The window consists of $N_{k}$ keyframes carefully selected from the latest images in every time interval based on co-visibility, and every keyframe contains detected Shi-Tomasi corners [shi1994good] tracked by the KLT sparse optical flow [lucas1981iterative]. We triangulate a corner with the estimated keyframe poses and camera intrinsics once it is consecutively tracked in $N_{t}$ keyframes. The newly triangulated SFM points will be combined with LiDAR points for supplementation.
@@ III-C Photo-Realistic Mapping using Gaussian Splatting
The mapping module receives sequential frames from the tracking module to perform reconstruction, each of which contains an estimated camera pose, an undistorted image, and colored LiDAR points combined with visual SFM points.
@@ III-C1 Initialization of 3D Gaussians
The photo-realistic Gaussian map is bootstrapped with all 3D points of the first received frame. To be specific, for each point, we initialize a new Gaussian centered at its position with the zeroth degree of $\mathbf{SH}$ filled with its RGB color, opacity set to be $o_{a}$, and rotation set to be the identity matrix. To mitigate the aliasing artifacts [yan2024multi], we assign smaller scales for the closer Gaussians, while larger scales for those far away from the image plane, namely $\frac{d}{f}\mathbf{e}$, where $\mathbf{e}$ is a $3\times 1$ vector filled with 1, while $d$ and $f$ denote the depth of the point in the current camera frame and the camera focal length, respectively.
@@ III-C3 Optimization
where $\bar{\mathbf{C}}$ and $\mathbf{C}$ are the ground-truth image and the image rendered by Eq. (5), respectively. $\mathcal{L}_{\mathrm{D}-\mathrm{SSIM}}$ is a D-SSIM term [kerbl20233d]. To handle the varying camera exposure, we additionally optimize a $3\times 4$ affine matrix $\mathbf{E}$ which is applied to the rendered color [kerbl2024hierarchical]. The first three columns are used for scaling, and the last column is used for offsetting.
@@ III-D Meticulous CUDA-related Designs for Acceleration
Implemented fully in C++ and CUDA, our SLAM system consists of a tracking module and a mapping module, running in parallel and exchanging data based on ROS. The mapping module is based on the LibTorch framework. Inspired by [mallick2024taming, radl2024stopthepop], we carefully integrate and design a series of strategies related to CUDA to boost the training of Gaussians, an important aspect overlooked in previous works.
@@ III-D1 Per-Gaussian Backpropagation
During the backward process in the original work [kerbl20233d], gradients are propagated from the pixels onto the Gaussians, accounting for the majority of the time consumption of the training. Specifically, each pixel corresponds to a different number of Gaussians, and the time cost of such a per-pixel backpropagation paradigm is limited to the pixel with the most Gaussians. All other CUDA threads have nothing to do but just rest until this pixel finishes backward. Also, the collisions of atomic gradient addition happen when the different pixels backpropagate to the same Gaussian. We instead perform per-Gaussian backpropagation and assign CUDA threads for all Gaussians in a tile, increasing the parallelism while decreasing collisions.
III-D2 Culling with Load Balancing
The vanilla 3DGS approximates the 2D splatted Gaussians as circles to determine relevant tiles. However, it is unfriendly for slim Gaussians which often appear in incremental mapping systems such as SLAM, since they contribute little to most of the tiles. Hence, we efficiently cull those unimportant tiles via load balance.
III-D3 Other Useful Tricks
For faster training of the Gaussian map representation, we further make use of the sparse Adam to only update Gaussians in the current camera frustum. We remove the background color in the forward and backward process since it is kept as 0 all the time. We directly pass the ground truth image into the rasterizer to construct the loss, thereby avoiding the need to launch an additional CUDA kernel. In addition, we temporarily store repeated divisions for subsequent reuse.

### v3

@@ I INTRODUCTION
Simultaneous localization and mapping (SLAM) is a pivotal technique in mixed reality and robotics. In recent years, emerging radiance field technologies such as Neural Radiance Fields (NeRF) [1] and 3D Gaussian Splatting (3DGS) [2] have carved out a new direction in the realm of SLAM, namely radiance-field-based SLAM [3], aiming to simultaneously localize accurately and construct photo-realistic 3D maps. This holds significant importance for enhancing the capabilities of robots to model, understand, and interact with natural environments in real time [4, 5, 6, 7].
Compared to NeRF which involves computationally intensive volumetric rendering based on ray tracing, 3DGS offers fast rendering speed with superior visual quality, showcasing a greater potential in real-time SLAM applications. Yet there still remain several issues for 3DGS-based SLAM. First, similar to existing NeRF-based SLAM [8, 9, 10, 11, 12, 13, 14, 15, 16], most 3DGS-based SLAM [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] focus mainly on indoor scenarios using RGB-D or RGB sensors. When it comes to unbounded outdoor scenes or encountering challenging conditions such as violent ego-motion, varying illumination, and lack of visual textures, the performance of existing 3DGS-based SLAM might dramatically diminish. Second, for the outdoor scenarios, the modeling of the sky and varying camera exposure becomes important. Ignoring them can incur significant artifacts to 3DGS in unbounded scenes. Third, real-time performance is still a bottleneck. Constraining the number of or the dimension of attributes of 3D Gaussian representation [18, 22] could be effective indoors but might lead to deteriorated performance in unbounded outdoor scenarios.
@@ II-A Photo-Realistic RGB-D or RGB SLAM
Given sequential RGB-D or RGB inputs, abundant works attain brilliant results of localization and photo-realistic mapping in indoor environments. iMAP [8], the first NeRF-based SLAM that employs the implicit neural representation to achieve watertight online reconstruction, has pioneered a new era in SLAM. As a follow-up, NICE-SLAM [9] combines MLPs with hierarchical feature grids, markedly improving the quality of scene representation and realizing outstanding performance in larger indoor rooms. Further, Vox-Fusion [10] utilizes octree to dynamically expand the volumetric neural implicit map, eliminating the need for pre-allocated grids. Adopting hash-grids, tri-planes, and neural point clouds as implicit neural representations respectively, Co-SLAM [11], ESLAM [12], and Point-SLAM [13] get enhancement in both localization and reconstruction. Meanwhile, HERO-SLAM [14] improves the robustness in the optimization of the neural map. NeRF employs time-consuming ray-based volume rendering, while 3DGS utilizes fast point-based rasterization, accelerating image view synthesis and producing promising rendering quality. Notably, GS-SLAM [17], SplaTAM [18], and Gaussian-SLAM [19] detailedly elucidate the significant advantages of 3DGS over existing map representations in SLAM tasks for online photo-realistic mapping. SGS-SLAM [27], SemGauss-SLAM [28], and NEDS-SLAM [29] integrate semantic feature embedding into Gaussians and demonstrate superb capabilities in online dense semantic mapping. Fusing Generalized Iterative Closest Point (G-ICP) and 3DGS, GS-ICP-SLAM [20] shares covariances of Gaussians between tracking and mapping to minimize redundant computations. CG-SLAM [21] additionally incorporates a depth uncertainty model to select valuable Gaussians and thereby improves tracking performance. By forcing each Gaussian to be either opaque or nearly transparent, RTG-SLAM [22] achieves real-time performance indoors with compact scene representation.
Several works have tried to operate solely on monocular images. Among them, NeRF-SLAM [15] and IG-SLAM [23] utilize the dense depth maps estimated from the tracking front-end DROID-SLAM [30] as additional information to supervise the training of Instant-NGP [31] and 3DGS. Correspondingly, NeRF-VO [16] and MGS-SLAM [24] instead employ the sparse visual odometry DPVO [32] as a faster front-end with network-predicted dense depth for supervision. Concurrently, Photo-SLAM [25] utilizes the classical visual odometry ORB-SLAM3 [33] for accurate pose estimation and reconstructs a hybrid Gaussian map with ORB features. MonoGS [26] introduces geometric regularization to address ambiguities in incremental reconstruction.
@@ II-B Photo-Realistic Multimodal SLAM
RGB-D and RGB approaches can realize fancy results in well-lit indoor scenes, but they would find it difficult to scale up to complex unbounded outdoor environments or cope with violent motions, lighting changes, and texture missing. Also, RGB-D sensors are often unavailable outdoors. To this end, fusing multimodal sensors such as LiDAR and IMU will be necessary and effective, which has been extensively proven in classical SLAM methods [34, 35, 36, 37, 38, 39, 40] characterized by robustness and accuracy. Nevertheless, their primary focus lies in geometric mapping, resulting in geometrically precise yet visually simple representations of the scene. URF [41], DrivingGaussian [42], StreetGaussians [43], PVG [44], LIV-GaussMap [45], and LetsGo [46] perform enhanced photorealistic reconstruction based on LiDAR-camera fusion, but all of them are offline methods taking a batch input.
Oriented to SLAM tasks, MM-Gaussian [47] leverages accurate geometric structure information from a solid-state LiDAR to substitute the depth of the RGB-D, thus scaling up to the outdoor scenes. Specifically, it first performs a pure LiDAR odometry [48] for pose estimations and further refines the pose via color and depth rendering loss. Similar to SplaTAM [18], it optimizes isotropic Gaussians without view-dependent effect initialized from LiDAR points. MM3DGS-SLAM [49] fuses RGB-D camera and IMU measurements in a loosely-coupled manner, to enable more accurate and scale-aware pose estimation. Both of the aforementioned multimodal SLAM approaches are capable of functioning in outdoor environments. However, their mapping performance is suboptimal due to the lack of sky modeling and the challenges posed by varying camera exposure in outdoor settings. Additionally, they are not specially optimized for real-time applications.
@@ III-A Preliminary: 3D Gaussian Splatting
To reconstruct a richly detailed dense map for fast and high-quality rendering, we represent the scene as a collection of anisotropic 3D Gaussians [2], each of which contains position $\bm{\mu}\in\mathbb{R}^{3}$ in the world coordinate, scale $\mathbf{S}\in\mathbb{R}^{3}$, rotation $\mathbf{R}\in\mathbb{R}^{3\times 3}$, opacity $o\in\mathbb{R}$, and three-degree spherical harmonics $\mathbf{SH}\in\mathbb{R}^{48}$ capturing the view-dependent appearance of the scene. Representing the Gaussian’s ellipsoidal shape, the covariance of each Gaussian is parameterized as:
@@
where $\mathbf{e}_{i}$ is a $3\times 1$ vector with its $i$-th element to be 1 and the other elements to be 0. Then $^{\mathbf{e}_{3}^{\top}\hat{\bm{\mu}}}$ is the depth $d$ of the Gaussian in the camera frame. $\pi_{c}(\cdot)$ denotes the projection operation which transforms a point on the normalized image plane to a pixel. $\mathbf{J}$ is the Jacobian of the affine approximation of the projective transformation and $\cdot]_{2\times 2}$ skips the third row and column to acquire a $2\times 2$ matrix. The projected 2D Gaussian affects the pixel $\bm{\rho}=\begin{bmatrix}u&v\end{bmatrix}^{\top}$ by the weight:
@@ III-B Multimodal Odometry and Mapping Data Preparation
Given data from multimodal sensors from an RGB camera, a 3D LiDAR, and an IMU, we adopt our previous work Coco-LIC [40], a continuous-time odometry to tightly fuse the sensor data, delivering robust, real-time, and accurate pose estimation even in challenging scenarios. It incrementally optimizes the trajectory for every 0.1 seconds via factor graph optimization, incorporating point-to-map LiDAR factors, frame-to-map visual factors, and inertial factors.
For photo-realistic reconstruction, 3D Gaussians in the map need to be bootstrapped by a set of 3D points. The better initialization leads to faster convergence and higher-quality rendering. Exactly, the LiDAR points provide handy and more precise geometric priors for Gaussian initialization [42]. Our system supports both solid-state and mechanical spinning LiDARs. Specifically, after finishing the sliding-window optimization of the trajectory segment in the latest 0.1 seconds, we first downsample the LiDAR points in this time interval by retaining one out of every $N_{l}$ points randomly. Then we colorize the remained LiDAR points by projection onto the latest image and remove those outside the corresponding camera FoV (field of view). Afterward, these colorized LiDAR points in the world frame are sent to the mapping thread, together with the posed latest image, regarded as a frame.
Though offering accurate geometric priors, the region scanned by the LiDAR is limited compared to the camera, especially for the repetitive mechanical spinning LiDAR. Thus, we propose maintaining a visual sliding window similar to VINS-Mono [50] to compute triangulated SFM points for compensation. The window consists of $N_{k}$ keyframes carefully selected from the latest images based on co-visibility, and every keyframe contains detected Shi-Tomasi corners [51] tracked by the KLT sparse optical flow [52]. We triangulate a corner with the estimated keyframe poses and camera intrinsics once it is consecutively tracked in $N_{t}$ keyframes. The newly triangulated SFM points will be combined with LiDAR points for supplementation.
@@ III-C Photo-Realistic Mapping using Gaussian Splatting
The mapping module receives sequential frames from the tracking module to perform reconstruction. Each frame consists of an estimated camera pose, an undistorted image, and colored LiDAR points combined with visual SFM points.
@@ III-C1 Initialization of 3D Gaussians
The photo-realistic Gaussian map is bootstrapped with all 3D points of the first received frame. To be specific, for each point, we initialize a new Gaussian centered at its position with the zeroth degree of $\mathbf{SH}$ filled with its RGB color, opacity set to be $o_{a}$, and rotation set to be the identity matrix. To mitigate the aliasing artifacts [53], we assign smaller scales for the closer Gaussians, while larger scales for those far away from the image plane, namely $\frac{d}{f}\mathbf{e}$, where $\mathbf{e}$ is a $3\times 1$ vector filled with 1, while $d$ and $f$ denote the depth of the point in the current camera frame and the camera focal length, respectively.
@@ III-C3 Optimization
where $\bar{\mathbf{C}}$ and $\mathbf{C}$ are the ground-truth image and the image rendered by Eq. (5), respectively. $\mathcal{L}_{\mathrm{D}-\mathrm{SSIM}}$ is a D-SSIM term [2]. To handle the varying camera exposure, we additionally optimize a $3\times 4$ affine matrix $\mathbf{E}$ which is applied to the rendered color [54]. The first three columns of $\mathbf{E}$ are used for scaling, and the last column is used for offsetting.
@@ III-D Meticulous CUDA-related Designs for Acceleration
Implemented fully in C++ and CUDA, our SLAM system consists of a tracking module and a mapping module, running in parallel and exchanging data based on ROS. The mapping module is based on the LibTorch framework. Inspired by [55, 56], we carefully integrate and design a series of strategies related to CUDA to boost the training of Gaussians, an important aspect overlooked in previous works.
@@ III-D1 Per-Gaussian Backpropagation
During the backward pass of Gaussian Splatting [2], gradients are propagated from pixels to Gaussians, constituting the most computationally expensive part of training. Specifically, each pixel corresponds to a varying number of Gaussians, and the overall time cost is dictated by the pixel with the highest number of Gaussians. Consequently, all other CUDA threads remain idle, waiting for this bottleneck pixel to complete its backpropagation. Additionally, atomic gradient addition conflicts arise when multiple pixels backpropagate to the same Gaussian. To mitigate this, we adopt a per-Gaussian backpropagation strategy [55], assigning CUDA threads to all Gaussians within a tile. This allows to enhance parallelism while reducing collisions.
III-D2 Tile Culling and Additional Tricks
The vanilla 3DGS [2] approximates projected 2D splatted Gaussians as circles to determine relevant tiles. However, this approach is suboptimal for slim 2D Gaussians, which frequently appear in incremental mapping systems such as SLAM, as they contribute minimally to most tiles selected by a circular approximation. In our implementation, we efficiently cull these weakly affected tiles through load balancing [56], enhancing computational efficiency. For faster training of the Gaussian map representation, we further make use of the sparse Adam [55] to only update Gaussians in the current camera frustum instead of optimizing all the Gaussians.